OpenStudy (anonymous):

Adults' tickets for the art museum cost $1.50 more than children's tickets. One day 200 adults' tickets and 160 children's tickets were sold at the museum. The income from the adults' tickets was $30 less than twice the income from the children's tickets. How much does each type of ticket cost?? Thanks for the help! :D

OpenStudy (texaschic101):

A = C + 1.50 A = 2C - 30 2C - 30 = C + 1.50 2C - C = 1.50 + 30 C = 31.50 A = 2(31.50) - 30 A = 63 - 30 A = 33 Now lets check... A = 2C - 30 33 = 2(31.50) - 30 33 = 63 - 30 33 = 33 (correct) Adult tickets (A) cost $ 33 Child tickets (C) cost $31.50